# Harmonic functions on Tutte’s embeddings and linearized Monge–Ampère equation

### v1

@@ 1.4. Convergence to harmonic functions in a non-trivial metric and t-surfaces in ${\mathbb{R}}^{2,2}$
    There exists a coordinate change $w=w(\zeta)$ such that we have ${\mathcal{L}}h=0$ if and only if $h$ is harmonic in $\zeta$. If this holds, then $\zeta$ is a conformal coordinate on $\Theta$.
@@
    The two differential equations ${\mathcal{L}}_{\varphi}h=0$ and ${\mathcal{L}}^{\times}_{\varphi}h=0$ are equivalent.
@@ 1.5. Organization of the paper
We recall the combinatorial construction of t-embeddings and t-surfaces associated with Tutte’s embedding and the modern discrete complex analysis framework developed in [CLR1] in Section 2. Section 3 is devoted to basic properties of elliptic equations in 2d. We prove main convergence theorems in Section 4 basing upon classical Dirichlet energy and Caccioppoli estimates. Section 5 contains the proof of the equivalence of properties (CONV)/(LIP) and property (RW). Theorem 4 is proved in Section 6.
@@ 2. T-surface of a harmonic embedding
Given a weighted planar graph $(\Gamma_{\delta},c)$ embedded into the complex plane harmonically, there are two discrete surfaces that can be associated with it. The first is the graph of the convex potential $\Phi_{\delta}$ associated with the embedding and discussed in the introduction; this surface belongs to the Euclidean space ${\mathbb{R}}^{3}$. The second is called a t-surface: this is a polygonal surface $\Theta_{\delta}$ embedded into ${\mathbb{R}}^{2,2}\cong{\mathbb{C}}^{1,1}$, which is obtained as the lift to this Minkowski space of the t-embedding associated with the harmonic embedding as described in [CLR1, Section 8.1]. The structure of $\Theta_{\delta}$ plays a crucial role in the discrete complex analysis technique developed in [CLR1] upon which our approach is based. In this section we focus on constructing this t-surface and reviewing basic facts about it. For shortness, we drop the index $\delta$ until the end of this section.
@@ 2.1. Definition of the t-surface
Note that there is a bijection between the edges of $\mathcal{V}$} and the edges of $\Gamma\cup\Gamma^{\ast}$. Also, there is a bijection between faces of $\mathcal{V}$} and all vertices of $\Gamma\cup\Gamma^{\ast}$ except the black boundary ones. The latter correspondence allows us to color the faces of $\mathcal{V}$} in black and white accordingly to the coloring of vertices of $\Gamma\cup\Gamma^{\ast}$; see Figure 2 for an example. We denote by ${\mathrm{B}}({\mathcal{V}})$ and ${\mathrm{W}}({\mathcal{V}})$ the sets of black and white faces of $\mathcal{V}$}, respectively. The next definition follows [CLR1, Section 8.1]:
@@
Following the terminology of [CLR1], each t-embedding $\mathcal{T}$} has an origami map $\mathcal{O}$} associated with it. Informally speaking, the mapping $\mathcal{O}$}, defined up to rotations and translations, is obtained from the map $\mathcal{T}$} by folding the plane along all the edges of $\mathcal{T}$} consequently (one can prove that this procedure is consistent). Note that the images of a given face of $\mathcal{V}$} in $\mathcal{T}$} and in $\mathcal{O}$} are isometric to each other, and that these images have the same orientation for white faces and the opposite ones for black faces. In the setup of this paper one can use an explicit definition
@@
see also equation (2.9) below and [CLR1, Section 8.1] for more details.
@@
see also equation (2.9) below. Recall that the function $\Phi$, defined by (1.3), is the convex piecewise linear potential associated with $\Gamma$ and that $\Psi=2\partial_{\bar{w}}\Phi=\partial_{x}\Phi+i\partial_{y}\Phi$.
@@
This follows directly from the definition of $\mathcal{T}$} and $\mathcal{O}$}; see also [CLR1, Section 8.1]. ∎
@@
According to Lemma 2.3 we can view $\mathcal{T}$} and $\mathcal{O}$} as pullbacks of the functions $w\mapsto\tfrac{1}{2}(w+\Psi)$ and $w\mapsto\tfrac{1}{2}(\overline{\Psi(w)-w})$ from the plane into which $\Gamma$ is embedded by $\mathcal{H}$} to the corresponding t-surface $\Theta$. (Note that, contrary to $\Psi$, the maps $\mathcal{T}$} and $\mathcal{O}$} are well-defined and continuous on $\Theta$.) If the potential $\Phi$ was smooth and uniformly convex, then one could easily prove that $\mathcal{O}$} is, locally, a $\kappa$-Lipschitz function of $\mathcal{T}$} for some $\kappa<1$ that depends only on the constant $\lambda$ in (1.4) or, equivalently, on the constant $\varkappa$ in (1.5); see also Lemma 5.1. Moreover, $\Phi$ is uniformly convex if and only if this is true. In our setup $\mathcal{O}$} is locally a linear function of $\mathcal{T}$}, which means that this $\kappa$-Lipschitzness property cannot hold at very small distances. However, it is reasonable to expect that property (CONV)/(LIP) implies that $\mathcal{O}$} is a $\kappa$-Lipschitz function of $\mathcal{T}$} starting from the scale $\delta$. The following definition appeared in [CLR1]:
@@
As explained in [CLR1, Section 8.1], $\Gamma$ and $\Gamma^{\ast}$ are T-graphs obtained from $\Theta=({\mathcal{T}},{\mathcal{O}})$ as projections ${\mathcal{T}}-\overline{{\mathcal{O}}}$ and ${\mathcal{T}}+\overline{{\mathcal{O}}}$ (cf. Lemma 2.3). The ellipticity estimate (2.7) for the random walk on such T-graphs then follows from [CLR1, Proposition 6.4]. To make this entirely clear, note that since (2.7) is local and $\Theta$ satisfies $\mathrm{Lip}(\kappa,\delta)$, it is enough to assume that the projection $\Theta\mapsto{\mathcal{T}}$ is one-to-one (cf. the proof of Lemma 2.6), that is, $\mathcal{T}$} is a proper t-embedding as defined in [CLR1]; then $\mathcal{O}$} is a concrete instance of its origami map. If we switch colors of faces of $\mathcal{T}$}, then $\pm\overline{{\mathcal{O}}}$ will represent the origami of the obtained t-embedding (see [CLR1, Section 2.2]), to which we can apply [CLR1, Proposition 6.4] to get the desired estimates for ${\mathcal{T}}-\overline{{\mathcal{O}}}$ and ${\mathcal{T}}+\overline{{\mathcal{O}}}$. ∎
@@
Instead of referring to [CLR1, Proposition 6.4] one can give a self-contained proof of Proposition 2.7 using the notion of discrete extremal length; see Remark 4.3 for further comments.
The following lemmas are standard corollaries of Proposition 2.7; see [CLR1, Sections 6.3, 6.4]):
@@
The claim now follows from [CLR1, Eq. (6.1) and Lemma 6.3]. ∎
@@ 2.4. T-holomorphic functions
Let $\Gamma$ be a weighted planar graph and $\Gamma^{\ast}$ be its dual. There is a standard way to define a discrete holomorphic function on the superposition of $\Gamma$ and $\Gamma^{\ast}$ by declaring it to be a pair of $H:\Gamma\to{\mathbb{R}}$ and $H^{\ast}:\Gamma^{\ast}\to{\mathbb{R}}$ that are harmonic conjugate to each other. Assume now that $\Gamma$ is embedded harmonically and $\mathcal{T}$} is the corresponding t-embedding; see Section 2.1. In this case the aforementioned discrete holomorphic functions correspond to t-white-holomorphic functions on $\mathcal{T}$} as defined in [CLR1]. It is convenient to introduce an alternative definition of the latter using the terminology developed in Section 2.1.
We begin by fixing an instance of the origami square root function $\eta$ associated with the t-embedding $\mathcal{T}$} and the origami map $\mathcal{O}$}; see [CLR1, Section 2.2]. The function $\eta$ is defined on the set ${\mathrm{B}}({\mathcal{V}})\cup{\mathrm{W}}({\mathcal{V}})$ of faces of $\mathcal{T}$}. Recall that $\mathcal{V}$} is the corner graph and there is a correspondence between ${\mathrm{B}}({\mathcal{V}})$ on the one side and vertices of $\Gamma$ and those of $\Gamma^{\ast}$ on the other, as well as between ${\mathrm{W}}({\mathcal{V}})$ and edges of $\Gamma$, which are also in a one-to-one correspondence with edges of $\Gamma^{\ast}$. (See Definition 2.1 and the discussion below it.) Define the function $\eta$ as follows:
@@
since ${\mathcal{H}}^{\ast}(v_{2}^{\ast})-{\mathcal{H}}^{\ast}(v_{1}^{\ast})=ic_{v_{1}v_{2}}({\mathcal{H}}(v_{2})-{\mathcal{H}}(v_{1}))$ by definition of the dual harmonic embedding. Taking into account that $F(b)$ is purely real when $b\in\Gamma$ and purely imaginary when $b\in\Gamma^{\ast}$ we conclude that the functions $H=F|_{\Gamma}$ and $H^{\ast}=-iF|_{\Gamma^{\ast}}$ are harmonic conjugate to each other. Thus, we obtained the definition of discrete holomorphicity that we have started with. We address the reader to [CLR1, Section 8.1] for more comments.
@@
In order to define ‘true complex values’ $G^{\bullet}$ of $G$, we triangulate each black face of $\Theta$, extend $H$ and $H^{\ast}$ linearly to thus obtained triangles $b\in{\mathrm{B}}_{\mathrm{spl}}({\mathcal{V}})$, and define $G^{\bullet}(b)$ to be the gradient of $H$ (if $b$ is a part of a face of $\Gamma^{\ast}$) or gradient of $H^{\ast}$ (if $b$ is a part of a face of $\Gamma$), viewed as a complex number and multiplied by $i$. Similarly to Lemma 2.14, this allows one to introduce a closed differential 1-form $G^{\bullet}d{\mathcal{T}}+\overline{G^{\bullet}}d{\mathcal{O}}$ on $\Theta$. We address the reader to [CLR1, Section 8.1] for more details.
@@ 3. Elliptic equations and closed 1-forms in 2D
In this section we collect some facts about linearized Monge–Ampère equation that we will need later. This is a classical topic at the crossroad of the theory of elliptic PDE and the optimal transport problem. We do not attempt to summarize the literature and only mention the foundational paper [caffarelli-gutierrez-1997-lin-MA] and the monograph [figalli-2017-MA] where the link with the transport problem is addressed. For our purposes it is enough to restrict the discussion to linearized Monge–Ampère equations with uniformly convex potentials. Such equations are uniformly elliptic which makes the standard theory of elliptic PDE applicable. To this end, we address the reader to [kenig2000new] and references therein for the standard properties of elliptic operators in divergence form with measurable coefficients, and the book [AstalaBook] as a standard reference to the interplay between Beltrami and elliptic equations in 2D.
@@
Classically [AstalaBook, Chapter 16.1.3], with every solution $h$ of the equation ${\mathcal{L}}h=0$ we can associate its $A$-harmonic conjugate $h^{\ast}$ defined by
@@
To this end, note that the difference $h-h_{0}$ is continuous in $\overline{B}$, has zero boundary values on $\partial B$ and satisfies (3.7) for each twice differentiable compactly supported test function $\xi$ in $B$. Now let $\xi$ be the (strong) solution of the nonhomogeneous equation ${\mathcal{L}}_{\varphi}\xi=\phi$ in $B$ with zero boundary values. As $\phi\in{\mathrm{C}}_{0}^{\infty}(B)$, classical Schauder estimates [Gilbarg-Trudinger, Chapter 6] imply that $\xi$ is ${\mathrm{C}}^{2}$-smooth inside $B$ and has uniformly bounded gradient in $B$ including near $\partial B$. Finally, for sufficiently small $\varepsilon>0$ denote $\xi_{\varepsilon}=\eta_{\varepsilon}\xi$, where $\eta_{\varepsilon}:B\to[0,1]$ is a smooth function such that $\eta_{\varepsilon}=1$ in $B(p,(1-2\varepsilon)r)$, $\eta_{\varepsilon}=0$ outside $B(p,(1-\varepsilon)r)$, and $|D^{2}\eta_{\varepsilon}|=O(\varepsilon^{-2})$. Since $\xi_{\varepsilon}\in{\mathrm{C}}_{0}^{2}(B)$, by applying (3.7) to $h-h_{0}$ and $\xi_{\varepsilon}$ we see that
@@
Classically (see [littman-stampacchia-weinberger-1963-regular-points, Section 6] or [taylor-kim-brown-2013-green-function] and references therein), the operator ${\mathcal{L}}_{\varphi}$ admits a Green’s function in a domain $\Omega$, that is, a function $G(w_{1},w_{2})$ that solves the equation
@@
(where $\delta_{w_{2}}$ is the delta measure at $w_{2}$) for each $w_{2}\in\Omega$ and vanishes along the boundary of $\Omega$. The standard theory of elliptic equations imply that $G$ is symmetric, positive on $\Omega\times\Omega\smallsetminus\mathrm{diag}$, and continuous on $\overline{\Omega}\times\overline{\Omega}\smallsetminus\mathrm{diag}$. Moreover, any non-negative continuous function $\widetilde{G}(\cdot,w_{2})$ in $\overline{\Omega}\smallsetminus\{w_{2}\}$ that vanishes at the boundary of $\Omega$ and satisfies the equation ${\mathcal{L}}_{\varphi}\widetilde{G}(\cdot,w_{2})=0$ coincides with $G(\cdot,w_{2})$ up to a multiplicative constant. Indeed, standard application of Harnack inequality (see e.g. [littman-stampacchia-weinberger-1963-regular-points, Section 7]) implies that such $\widetilde{G}(w_{1},w_{2})$ has the same growth as $G(w_{1},w_{2})$ when $w_{1}\to w_{2}$, which implies the assertion due to a classical result from [serrin-weinberger-1966-isolated-singularities]. The multiplicative constant can be fixed by computing the monodromy of the $A$-harmonic conjugate:
@@
By the standard continuity arguments (e.g., see [littman-stampacchia-weinberger-1963-regular-points, Section 5]) it is enough to assume that $\varphi$ is smooth, thus, by Schauder estimates, $G$ is smooth as well. Fix a smooth function $\xi:{\mathbb{R}}\to{\mathbb{R}}$ with a small support that is constant in a vicinity of 0. Let $m$ be the monodromy of $G^{\ast}(\cdot,w_{2})$. Then, using the definition of $A$-harmonic conjugate and the definition of the Hodge star $\ast$, and integrating along concentric circles around $w_{2}$ we can compute

### v2

@@ 1.4. Convergence to harmonic functions in a non-trivial metric and t-surfaces in ${\mathbb{R}}^{2,2}$
    There exists a coordinate change $w=w(\zeta)$ such that we have ${\mathcal{L}}_{\varphi}h=0$ if and only if $h$ is harmonic in $\zeta$. If this holds, then $\zeta$ is a conformal coordinate on $\Theta$.
@@
    The two differential equations ${\mathcal{L}}_{\varphi}h=0$ and ${\mathcal{L}}^{\times}_{\varphi}h=0$ are equivalent (have the same space of solutions on any open set $U^{\prime}\subset U$).
@@ 1.5. Organization of the paper
We recall the combinatorial construction of t-embeddings and t-surfaces associated with Tutte’s embedding and the modern discrete complex analysis framework developed in [11] in Section 2. Section 3 is devoted to basic properties of elliptic equations in 2d. We prove main convergence theorems in Section 4 basing upon classical Dirichlet energy and Caccioppoli estimates. Section 5 contains the proof of the equivalence of properties (CONV)/(LIP) and property (RW). Theorem 4 is proved in Section 6.
@@ 2. T-surface of a harmonic embedding
Given a weighted planar graph $(\Gamma_{\delta},c)$ embedded into the complex plane harmonically, there are two discrete surfaces that can be associated with it. The first is the graph of the Maxwell–Cremona potential $\Phi_{\delta}$ associated with the embedding and discussed in the introduction; this surface belongs to the Euclidean space ${\mathbb{R}}^{3}$. The second is called a t-surface: this is a polygonal surface $\Theta_{\delta}$ embedded into ${\mathbb{R}}^{2,2}\cong{\mathbb{C}}^{1,1}$, which is obtained as the lift to this Minkowski space of the t-embedding associated with the harmonic embedding as described in [11, Section 8.1]. The structure of $\Theta_{\delta}$ plays a crucial role in the discrete complex analysis technique developed in [11] upon which our approach is based. In this section we focus on constructing this t-surface and reviewing basic facts about it. For shortness, we drop the index $\delta$ until the end of this section.
@@ 2.1. Definition of the t-surface
Note that there is a bijection between the edges of $\mathcal{V}$} and the edges of $\Gamma\cup\Gamma^{\ast}$. Also, there is a bijection between faces of $\mathcal{V}$} and all vertices of $\Gamma\cup\Gamma^{\ast}$ except the black boundary ones. The latter correspondence allows us to color the faces of $\mathcal{V}$} in black and white accordingly to the coloring of vertices of $\Gamma\cup\Gamma^{\ast}$; see Figure 2 for an example. We denote by ${\mathrm{B}}({\mathcal{V}})$ and ${\mathrm{W}}({\mathcal{V}})$ the sets of black and white faces of $\mathcal{V}$}, respectively. The next definition follows [11, Section 8.1]:
@@
Following the terminology of [11], each t-embedding $\mathcal{T}$} has an origami map $\mathcal{O}$} associated with it. Informally speaking, the mapping $\mathcal{O}$}, defined up to rotations and translations, is obtained from the map $\mathcal{T}$} by folding the plane along all the edges of $\mathcal{T}$} consequently (one can prove that this procedure is consistent). Note that the images of a given face of $\mathcal{V}$} in $\mathcal{T}$} and in $\mathcal{O}$} are isometric to each other, and that these images have the same orientation for white faces and the opposite ones for black faces. In the setup of this paper one can use an explicit definition
@@
see also equation (2.9) below and [11, Section 8.1] for more details.
@@
see also equation (2.9) below. Recall that the function $\Phi$, defined by (1.3), is the Maxwell–Cremona potential associated with $\Gamma$ and that $\Psi=2\partial_{\bar{w}}\Phi=\partial_{x}\Phi+i\partial_{y}\Phi$.
@@
This follows directly from the definition of $\mathcal{T}$} and $\mathcal{O}$}; see also [11, Section 8.1]. ∎
@@
According to Lemma 2.3 we can view $\mathcal{T}$} and $\mathcal{O}$} as pullbacks of the functions $w\mapsto\tfrac{1}{2}(w+\Psi)$ and $w\mapsto\tfrac{1}{2}(\overline{\Psi(w)-w})$ from the plane into which $\Gamma$ is embedded by $\mathcal{H}$} to the corresponding t-surface $\Theta$. (Note that, contrary to $\Psi$, the maps $\mathcal{T}$} and $\mathcal{O}$} are well-defined and continuous on $\Theta$.) If the potential $\Phi$ was smooth and uniformly convex, then one could easily prove that $\mathcal{O}$} is, locally, a $\kappa$-Lipschitz function of $\mathcal{T}$} for some $\kappa<1$ that depends only on the constant $\lambda$ in (1.4) or, equivalently, on the constant $\varkappa$ in (1.5); see also Lemma 5.1. Moreover, $\Phi$ is uniformly convex if and only if this is true. In our setup $\mathcal{O}$} is locally a linear function of $\mathcal{T}$}, which means that this $\kappa$-Lipschitzness property cannot hold at very small distances. However, it is reasonable to expect that property (CONV)/(LIP) implies that $\mathcal{O}$} is a $\kappa$-Lipschitz function of $\mathcal{T}$} starting from the scale $\delta$. The following definition appeared in [11]:
@@
As explained in [11, Section 8.1], $\Gamma$ and $\Gamma^{\ast}$ are T-graphs obtained from $\Theta=({\mathcal{T}},{\mathcal{O}})$ as projections ${\mathcal{T}}-\overline{{\mathcal{O}}}$ and ${\mathcal{T}}+\overline{{\mathcal{O}}}$ (cf. Lemma 2.3). The ellipticity estimate (2.7) for the random walk on such T-graphs then follows from [11, Proposition 6.4]. To make this entirely clear, note that since (2.7) is local and $\Theta$ satisfies $\mathrm{Lip}(\kappa,\delta)$, it is enough to assume that the projection $\Theta\mapsto{\mathcal{T}}$ is one-to-one (cf. the proof of Lemma 2.6), that is, $\mathcal{T}$} is a proper t-embedding as defined in [11]; then $\mathcal{O}$} is a concrete instance of its origami map. If we switch colors of faces of $\mathcal{T}$}, then $\pm\overline{{\mathcal{O}}}$ will represent the origami of the obtained t-embedding (see [11, Section 2.2]), to which we can apply [11, Proposition 6.4] to get the desired estimates for ${\mathcal{T}}-\overline{{\mathcal{O}}}$ and ${\mathcal{T}}+\overline{{\mathcal{O}}}$. ∎
@@
Instead of referring to [11, Proposition 6.4] one can give a self-contained proof of Proposition 2.7 using the notion of discrete extremal length; see Remark 4.3 for further comments.
The following lemmas are standard corollaries of Proposition 2.7; see [11, Sections 6.3, 6.4]):
@@
The claim now follows from [11, Eq. (6.1) and Lemma 6.3]. ∎
@@ 2.4. T-holomorphic functions
Let $\Gamma$ be a weighted planar graph and $\Gamma^{\ast}$ be its dual. There is a standard way to define a discrete holomorphic function on the superposition of $\Gamma$ and $\Gamma^{\ast}$ by declaring it to be a pair of $H:\Gamma\to{\mathbb{R}}$ and $H^{\ast}:\Gamma^{\ast}\to{\mathbb{R}}$ that are harmonic conjugate to each other. Assume now that $\Gamma$ is embedded harmonically and $\mathcal{T}$} is the corresponding t-embedding; see Section 2.1. In this case the aforementioned discrete holomorphic functions correspond to t-white-holomorphic functions on $\mathcal{T}$} as defined in [11]. It is convenient to introduce an alternative definition of the latter using the terminology developed in Section 2.1.
We begin by fixing an instance of the origami square root function $\eta$ associated with the t-embedding $\mathcal{T}$} and the origami map $\mathcal{O}$}; see [11, Section 2.2]. The function $\eta$ is defined on the set ${\mathrm{B}}({\mathcal{V}})\cup{\mathrm{W}}({\mathcal{V}})$ of faces of $\mathcal{T}$}. Recall that $\mathcal{V}$} is the corner graph and there is a correspondence between ${\mathrm{B}}({\mathcal{V}})$ on the one side and vertices of $\Gamma$ and those of $\Gamma^{\ast}$ on the other, as well as between ${\mathrm{W}}({\mathcal{V}})$ and edges of $\Gamma$, which are also in a one-to-one correspondence with edges of $\Gamma^{\ast}$. (See Definition 2.1 and the discussion below it.) Define the function $\eta$ as follows:
@@
since ${\mathcal{H}}^{\ast}(v_{2}^{\ast})-{\mathcal{H}}^{\ast}(v_{1}^{\ast})=ic_{v_{1}v_{2}}({\mathcal{H}}(v_{2})-{\mathcal{H}}(v_{1}))$ by definition of the dual harmonic embedding. Taking into account that $F(b)$ is purely real when $b\in\Gamma$ and purely imaginary when $b\in\Gamma^{\ast}$ we conclude that the functions $H=F|_{\Gamma}$ and $H^{\ast}=-iF|_{\Gamma^{\ast}}$ are harmonic conjugate to each other. Thus, we obtained the definition of discrete holomorphicity that we have started with. We address the reader to [11, Section 8.1] for more comments.
@@
In order to define ‘true complex values’ $G^{\bullet}$ of $G$, we triangulate each black face of $\Theta$, extend $H$ and $H^{\ast}$ linearly to thus obtained triangles $b\in{\mathrm{B}}_{\mathrm{spl}}({\mathcal{V}})$, and define $G^{\bullet}(b)$ to be the gradient of $H$ (if $b$ is a part of a face of $\Gamma^{\ast}$) or gradient of $H^{\ast}$ (if $b$ is a part of a face of $\Gamma$), viewed as a complex number and multiplied by $i$. Similarly to Lemma 2.14, this allows one to introduce a closed differential 1-form $G^{\bullet}d{\mathcal{T}}+\overline{G^{\bullet}}d{\mathcal{O}}$ on $\Theta$. We address the reader to [11, Section 8.1] for more details.
@@ 3. Elliptic equations and closed 1-forms in 2D
In this section we collect some facts about linearized Monge–Ampère equation that we will need later. This is a classical topic at the crossroad of the theory of elliptic PDE and the optimal transport problem. We do not attempt to summarize the literature and only mention the foundational paper [9] and the monograph [20] where the link with the transport problem is addressed. For our purposes it is enough to restrict the discussion to linearized Monge–Ampère equations with uniformly convex potentials. Such equations are uniformly elliptic which makes the standard theory of elliptic PDE applicable. To this end, we address the reader to [27] and references therein for the standard properties of elliptic operators in divergence form with measurable coefficients, and the book [1] as a standard reference to the interplay between Beltrami and elliptic equations in 2D.
@@
Classically [1, Chapter 16.1.3], with every solution $h$ of the equation ${\mathcal{L}}h=0$ we can associate its $A$-harmonic conjugate $h^{\ast}$ defined by
@@
To this end, note that the difference $h-h_{0}$ is continuous in $\overline{B}$, has zero boundary values on $\partial B$ and satisfies (3.7) for each twice differentiable compactly supported test function $\xi$ in $B$. Now let $\xi$ be the (strong) solution of the nonhomogeneous equation ${\mathcal{L}}_{\varphi}\xi=\phi$ in $B$ with zero boundary values. As $\phi\in{\mathrm{C}}_{0}^{\infty}(B)$, classical Schauder estimates [22, Chapter 6] imply that $\xi$ is ${\mathrm{C}}^{2}$-smooth inside $B$ and has uniformly bounded gradient in $B$ including near $\partial B$. Finally, for sufficiently small $\varepsilon>0$ denote $\xi_{\varepsilon}=\eta_{\varepsilon}\xi$, where $\eta_{\varepsilon}:B\to[0,1]$ is a smooth function such that $\eta_{\varepsilon}=1$ in $B(p,(1-2\varepsilon)r)$, $\eta_{\varepsilon}=0$ outside $B(p,(1-\varepsilon)r)$, and $|D^{2}\eta_{\varepsilon}|=O(\varepsilon^{-2})$. Since $\xi_{\varepsilon}\in{\mathrm{C}}_{0}^{2}(B)$, by applying (3.7) to $h-h_{0}$ and $\xi_{\varepsilon}$ we see that
@@
Classically (see [31, Section 6] or [44] and references therein), the operator ${\mathcal{L}}_{\varphi}$ admits a Green’s function in a domain $\Omega$, that is, a function $G(w_{1},w_{2})$ that solves the equation
@@
(where $\delta_{w_{2}}$ is the delta measure at $w_{2}$) for each $w_{2}\in\Omega$ and vanishes along the boundary of $\Omega$. The standard theory of elliptic equations imply that $G$ is symmetric, positive on $\Omega\times\Omega\smallsetminus\mathrm{diag}$, and continuous on $\overline{\Omega}\times\overline{\Omega}\smallsetminus\mathrm{diag}$. Moreover, any non-negative continuous function $\widetilde{G}(\cdot,w_{2})$ in $\overline{\Omega}\smallsetminus\{w_{2}\}$ that vanishes at the boundary of $\Omega$ and satisfies the equation ${\mathcal{L}}_{\varphi}\widetilde{G}(\cdot,w_{2})=0$ coincides with $G(\cdot,w_{2})$ up to a multiplicative constant. Indeed, standard application of Harnack inequality (see e.g. [31, Section 7]) implies that such $\widetilde{G}(w_{1},w_{2})$ has the same growth as $G(w_{1},w_{2})$ when $w_{1}\to w_{2}$, which implies the assertion due to a classical result from [41]. The multiplicative constant can be fixed by computing the monodromy of the $A$-harmonic conjugate:
@@
By the standard continuity arguments (e.g., see [31, Section 5]) it is enough to assume that $\varphi$ is smooth, thus, by Schauder estimates, $G$ is smooth as well. Fix a smooth function $\xi:{\mathbb{R}}\to{\mathbb{R}}$ with a small support that is constant in a vicinity of 0. Let $m$ be the monodromy of $G^{\ast}(\cdot,w_{2})$. Then, using the definition of $A$-harmonic conjugate and the definition of the Hodge star $\ast$, and integrating along concentric circles around $w_{2}$ we can compute